How many gallons of a 60 percent acid solution must be mixed with 80 gallons of a 14 percent solution to obtain a solution that is 50 percent acid?
Answer by checkley71(8403)

.6x+.14*80=.5(80+x)
.6x+11.2=40+.5x
.6x-.5x=40-11.2
.1x=28.8
x=28.8/.1
x=288 gallons of 60 % acid is needed.
proof
.6*288+.14*80=.5(80+288)
172.8+11.2=.5*368
184=184